OpenStudy (anonymous):

differentiate (x-1/x^2)^1001

OpenStudy (anonymous):

\[\frac{\partial }{\partial x}\left(x-\frac{1}{x^2}\right)^{1001}=1001 \left(\frac{2}{x^3}+1\right) \left(x-\frac{1}{x^2}\right)^{1000} \]